# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ Appendix B Necessary and Sufficient Conditions for Maximum-Likelihood Decoding
Regarding the converse part, while we stated Lemma 3(a) for all $\ell\in\big{[}1,\frac{k}{\log k}\big{]}$ for consistency with part (b), it will turn out to be sufficient to restrict attention to $\ell=1$ with a single feasible $(C,\zeta)\in(0,\infty)\times(0,1)$. We thus specialize Lemma 3(a) to $\ell=1$ with a simpler set of notation. Specifically, if $\mathcal{J}=\{j\}$ for some $j\in s$, then we will simply write $\mathcal{M}_{\mathcal{J}},\mathcal{M}_{\mathcal{J}1},\mathcal{M}_{\mathcal{J}0}$ respectively as $\mathcal{M}_{j},\mathcal{M}_{j1},\mathcal{M}_{j0}$, and their cardinalities are denoted by $M_{j},M_{j1},M_{j0}$, respectively. When further given $\mathcal{J}^{\prime}=\{j^{\prime}\}$ for some $j^{\prime}\in[p]\setminus s$, we write $G_{\mathcal{J},\mathcal{J}^{\prime},1},G_{\mathcal{J},\mathcal{J}^{\prime},2},\mathcal{K}_{\ell,C,\zeta},k_{\ell,C,\zeta}$ in Lemma 3 as $G_{j,j^{\prime},1},G_{j,j^{\prime},2},\mathcal{K}_{C,\zeta},k_{C,\zeta}$, respectively. Then, Lemma 8(a) specializes to the following.
Corollary 1.
@@
$\mathcal{K}_{C,\zeta}=\Big{\{}j\in s:M_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}\Big{\}}$ (35)
@@
$G_{j,j^{\prime},1}-G_{j,j^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k},$ (36)
then $\mathcal{L}(s^{\prime})>\mathcal{L}(s)$ with $s^{\prime}=(s\setminus\{j\})\cup\{j^{\prime}\}$, which implies the failure of MLE (29) .
@@ Appendix C $-\ell$ Converse Analysis for Bernoulli Designs
Let $\mathcal{S}_{k}=\{S^{\prime}\subset[p]:|S^{\prime}|=k\}$ and recall that we consider the prior $S\sim\mathrm{Unif}(\mathcal{S}_{k})$. As noted in Appendix B, it suffices to prove the converse bound for the optimal decoder, that is, maximum likelihood estimation (MLE) given in (29). By symmetry with respect to re-ordering items, we can simply assume that the actual defective set is $s=[k]$, i.e., items $1,\dotsc,k$ are defective, and items $k+1,\dotsc,p$ are non-defective.
@@
For $j\in s$ and $j^{\prime}\in[p]\setminus s$ we will use the notation $\mathcal{M}_{j},\mathcal{M}_{j0},\mathcal{M}_{j1},\mathcal{N}_{0},\mathcal{N}_{00},\mathcal{N}_{01}$ and the corresponding cardinalities from Corollary 1. Note that these index sets and their cardinalities are deterministic given $\mathbf{X}_{s}$ and $\mathbf{Y}$. We consider $\ell=1$, and say that $(C,\zeta)\in(0,\infty)\times(0,1)$ is feasible if $\frac{Cn\nu e^{-\nu}}{k}$ and $\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}$ are integers. Given feasible $(C,\zeta)$ we will consider $\mathcal{K}_{C,\zeta}=\{j\in s:M_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}\}$ with cardinality $k_{C,\zeta}=|\mathcal{K}_{C,\zeta}|$ as in (35). For given $(j,j^{\prime})\in\mathcal{K}_{C,\zeta}\times([p]\setminus s)$, recall that $G_{j,j^{\prime},1}$ denotes the number of tests in $\mathcal{N}_{01}\cup\mathcal{M}_{j1}$ that contain the non-defective item $j^{\prime}$, and $G_{j,j^{\prime},2}$ denotes number of tests in $\mathcal{N}_{00}\cup\mathcal{M}_{j0}$ that contain the non-defective item $j^{\prime}$; see Corollary 1.
We first invoke Corollary 1 to identify the sufficient condition for the failure of MLE. In particular, MLE fails if for some feasible $(C,\zeta)\in(0,\infty)\times(0,1)$, the following two conditions (C1) and (C2) simultaneously hold:
@@
  • (C2) There exist $j\in\mathcal{K}_{C,\zeta}$ and $k+1\leq j^{\prime}\leq p$ such that
    $G_{j,j^{\prime},1}-G_{j,j^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}.$ (37)
@@
We first derive a condition for ensuring $k_{C,\zeta}\geq 1$ with $1-o(1)$ probability. Fix a constant $\xi<1$ (to be chosen close to $1$ later) and suppose $k^{\xi}$ is an integer for convenience. For a given $j\in s$ and any test (say, with index $i$), we define the probability of test $i$ containing only one defective $j$ as
$P_{1}=\mathbb{P}(i\in\mathcal{M}_{j})=\frac{\nu}{k}\Big{(}1-\frac{\nu}{k}\Big{)}^{k-1}=\frac{\nu e^{-\nu}(1+o(1))}{k}.$ (38)
@@
$\|\bm{S}^{\prime}-\bm{T}\|_{\mathrm{TV}}=\left(\sum_{j=1}^{k}p_{j}\right)\left(\frac{1}{\sqrt{2\pi e}}+O\left(\frac{1}{\sqrt{N\sum_{j=1}^{k}p_{j}}}\right)\right).$ (39)
By Lemma 4 and the fact that $n,k\to\infty$ as $p\to\infty$, we have
$d_{\mathrm{TV}}(\bm{M}_{\zeta},\hat{\bm{M}}_{\zeta})$ $=k^{\xi}P_{1}\Big{(}\frac{1}{\sqrt{2\pi e}}+O\Big{(}\frac{1}{\sqrt{nk^{\zeta}P_{1}}}\Big{)}\Big{)}$ (40)
@@
$k_{C,\zeta}$ $=|\mathcal{K}_{C,\zeta}|\geq|\mathcal{K}_{C,\zeta}\cap[k^{\xi}]|=\sum_{j=1}^{k^{\xi}}\mathbbm{1}\Big{(}M_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta Cn\nu e^{-\nu}}{k}\Big{)}.$ (41)
@@
$\mathbb{P}(k_{C,\zeta}=0)$ $\leq\mathbb{P}\left(\sum_{j=1}^{k^{\xi}}\mathbbm{1}\Big{(}M_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta Cn\nu e^{-\nu}}{k}\Big{)}=0\right)$ (42)
@@
$=\mathbb{P}\left(\sum_{j=1}^{k^{\xi}}\mathbbm{1}\Big{(}\hat{M}_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta Cn\nu e^{-\nu}}{k}\Big{)}=0\right)+o(1)$ (43)
$=\left[1-\mathbb{P}\left(\hat{M}_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta Cn\nu e^{-\nu}}{k}\right)\right]^{k^{\xi}}+o(1)$ (44)
$=\left[1-\mathbb{P}\left(\hat{M}_{j}=\frac{Cn\nu e^{-\nu}}{k}\right)\mathbb{P}\left(\text{Bin}\Big{(}\frac{Cn\nu e^{-\nu}}{k},\rho\Big{)}=\frac{\zeta Cn\nu e^{-\nu}}{k}\right)\right]^{k^{\xi}}+o(1),$ (45)
@@
where (44) holds because the $\hat{M}_{j}$ are independent, and the $\frac{Cn\nu e^{-\nu}}{k}$ tests in $\mathcal{M}_{j}$ are disjoint for different $j\in[k^{\xi}]$ (by definition of $\mathcal{M}_{j}$), so the $k^{\xi}$ summands in (43) are independent; then, (45) holds because $\hat{M}_{j}$ is only related to the randomness of $\mathbf{X}_{s}$, while given $M_{j}=\frac{Cn\nu e^{-\nu}}{k}$, $M_{j0}\sim\text{Bin}(\frac{Cn\nu e^{-\nu}}{k},\rho)$ is only related to the randomness of $\mathbf{Z}$. Since $\hat{M}_{j}\sim\text{Poi}(nP_{1})$, we have
$\mathbb{P}\left(\hat{M}_{j}=\frac{Cn\nu e^{-\nu}}{k}\right)$ $=e^{-nP_{1}}\frac{(nP_{1})^{\frac{Cn\nu e^{-\nu}}{k}}}{\big{(}\frac{Cn\nu e^{-\nu}}{k}\big{)}!}$ (46)
$=\frac{\exp\big{(}-\frac{n\nu e^{-\nu}(1+o(1))}{k}\big{)}}{(C(1+o(1)))^{\frac{Cn\nu e^{-\nu}}{k}}}\cdot\frac{\big{(}\frac{Cn\nu e^{-\nu}}{k}\big{)}^{\frac{Cn\nu e^{-\nu}}{k}}}{\big{(}\frac{Cn\nu e^{-\nu}}{k}\big{)}!}$ (47)
$\geq\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{(}C\log C+1+o(1)\big{)}\right)\cdot\exp\left(\frac{Cn\nu e^{-\nu}}{k}(1+o(1))\right)$ (48)
$=\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{(}C\log C-C+1+o(1)\big{)}\right),$ (49)
@@
where we substitute (38) in (46) and use Stirling’s approximation (specifically, we use $\frac{w^{w}}{w!}\geq\frac{e^{w-1}}{w}=e^{w(1+o(1))}$ for $w\to\infty$ in (48) with $w=\frac{Cn\nu e^{-\nu}}{k}$). Moreover, by anti-concentration ((26) in Lemma 1), we have
$\mathbb{P}\left(\text{Bin}\Big{(}\frac{Cn\nu e^{-\nu}}{k},\rho\Big{)}=\frac{\zeta Cn\nu e^{-\nu}}{k}\right)$ $\geq\frac{1}{\sqrt{\frac{2Cn\nu e^{-\nu}}{k}}}\exp\left(-\frac{Cn\nu e^{-\nu}}{k}D(\zeta\|\rho)\right)$ (50)
$=\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{(}C\cdot D(\zeta\|\rho)+o(1)\big{)}\right),$ (51)
@@
where (51) holds because $\frac{n}{k}=\Theta(\log k)\to\infty$ allows us to incorporate the leading factor into the $o(1)$ term in the exponent; analogous simplifications will be used in the subsequent analysis. Combining the bounds in (49) and (51), we obtain
$\mathbb{P}\left(\hat{M}_{j}=\frac{Cn\nu e^{-\nu}}{k}\right)\mathbb{P}\left(\text{Bin}\Big{(}\frac{Cn\nu e^{-\nu}}{k},\rho\Big{)}=\frac{\zeta Cn\nu e^{-\nu}}{k}\right)$ (52)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \geq\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{[}C\log C-C+C\cdot D(\zeta\|\rho)+1+o(1)\big{]}\right):=P_{2}.$ (53)
Substituting this into (45), we obtain
$\mathbb{P}(k_{C,\zeta}=0)$ $\leq(1-P_{2})^{k^{\xi}}+o(1)$ (54)
$\leq\exp\big{(}-k^{\xi}P_{2}\big{)}+o(1).$ (55)
@@
$f_{1}(C,\zeta,\rho):=C\log C-C+C\cdot D(\zeta\|\rho)+1,$ (56)
and noting that $\eta_{1}\log\frac{p}{k}\to\infty$ for any given $\eta_{1}>0$, the condition $k^{\xi}P_{2}\to\infty$ can be ensured by
$\xi\log k-\frac{n\nu e^{-\nu}}{k}\big{(}f_{1}(C,\zeta,\rho)+o(1)\big{)}\geq\eta_{1}\log\frac{p}{k}.$ (57)
Moreover, by $k=\Theta(p^{\theta})$ we have $\log k\sim\frac{\theta}{1-\theta}\log\frac{p}{k}$. Substituting this into (57), letting $\xi$ be arbitrarily close to $1$ and rearranging, we find that (C1) holds with $1-o(1)$ probability if
$n\leq\frac{(1-\eta_{1})\frac{\theta}{1-\theta}k\log\frac{p}{k}}{\nu e^{-\nu}\big{(}f_{1}(C,\zeta,\rho)+o(1)\big{)}}$ (58)
holds for some $\eta_{1}>0$ possibly different from the one in (57).
@@
Given $k_{C,\zeta}\geq 1$ and fixing an arbitrary index $j\in\mathcal{K}_{C,\zeta}$, in this part we identify a sufficient condition for (C2). Most of our analysis will be conditioned on $(\mathbf{X}_{s},\mathbf{Y})$ and only based on the randomness of $\mathbf{X}_{[p]\setminus s}$, but we first deduce the high-probability behaviour of the quantities $N_{01}=|\mathcal{N}_{01}|$ and $N_{00}=|\mathcal{N}_{00}|$ depending on randomness of $(\mathbf{X}_{s},\mathbf{Y})$. Specifically, $\mathcal{N}_{01}$ ($\mathcal{N}_{00}$, resp.) consists of the positive (negative, resp.) tests containing no defective, and since the probability of a specific test containing no defective is given by $(1-\frac{\nu}{k})^{k}=e^{-\nu}(1+o(1))$, we have $N_{01}\sim\text{Bin}(n,\rho e^{-\nu}(1+o(1)))$ and $N_{00}\sim\text{Bin}(n,(1-\rho)e^{-\nu}(1+o(1)))$. Then, by Hoeffding’s inequality, we have that
$\mathscr{A}_{1}:=\big{\{}N_{01}=\rho ne^{-\nu}(1+o(1)),\leavevmode\nobreak\ N_{00}=(1-\rho)ne^{-\nu}(1+o(1))\big{\}}$ (59)
@@
For the specific $j\in\mathcal{K}_{C,\zeta}$ we have $M_{j0}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}=o(n)$ and $M_{j1}=\frac{(1-\zeta)\cdot Cn\nu e^{-\nu}}{k}=o(n)$, and thus they are negligible compared to $N_{01},N_{00}$ given by $\mathscr{A}_{1}$ in (59):
$|\mathcal{N}_{01}\cup\mathcal{M}_{j1}|=N_{01}+M_{j1}=\rho ne^{-\nu}(1+o(1)),$ (60)
$|\mathcal{N}_{00}\cup\mathcal{M}_{j0}|=N_{00}+M_{j0}=(1-\rho)ne^{-\nu}(1+o(1)).$ (61)

### 23 Oct 2024 v2

@@ Appendix B Necessary and Sufficient Conditions for Maximum-Likelihood Decoding
Regarding the converse part, while we stated Lemma B.1(a) for all $\ell\in\big{[}1,\frac{k}{\log k}\big{]}$ for consistency with part (b), it will turn out to be sufficient to restrict attention to $\ell=1$ with a single feasible $(C,\zeta)\in(0,\infty)\times(0,1)$. We thus specialize Lemma B.1(a) to $\ell=1$ with a simpler set of notation. Specifically, if $\mathcal{J}=\{j\}$ for some $j\in s$, then we will simply write $\mathcal{M}_{\mathcal{J}},\mathcal{M}_{\mathcal{J}1},\mathcal{M}_{\mathcal{J}0}$ as $\mathcal{M}_{j},\mathcal{M}_{j1},\mathcal{M}_{j0}$ respectively, and their cardinalities are denoted by $M_{j},M_{j1},M_{j0}$ respectively. Similarly, when it further holds that $\mathcal{J}^{\prime}=\{j^{\prime}\}$ for some $j^{\prime}\in[p]\setminus s$, we write $G_{\mathcal{J},\mathcal{J}^{\prime},1},G_{\mathcal{J},\mathcal{J}^{\prime},2},\mathcal{K}_{\ell,C,\zeta},k_{\ell,C,\zeta}$ in Lemma B.1 as $G_{j,j^{\prime},1},G_{j,j^{\prime},2},\mathcal{K}_{C,\zeta},k_{C,\zeta}$, respectively. Then, Lemma 8(a) specializes to the following.
Corollary B.1.
@@
$\mathcal{K}_{C,\zeta}=\Big{\{}j\in s:M_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}\Big{\}}$ (B.7)
@@
$G_{j,j^{\prime},1}-G_{j,j^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k},$ (B.8)
then $\mathcal{L}(s^{\prime})>\mathcal{L}(s)$ with $s^{\prime}=(s\setminus\{j\})\cup\{j^{\prime}\}$, which implies the failure of the MLE decoder (B.1) .
@@ Appendix C $-\ell$ Converse Analysis for Bernoulli Designs
Let $\mathcal{S}_{k}=\{S^{\prime}\subset[p]:|S^{\prime}|=k\}$ and recall that we consider the prior $S\sim\mathrm{Unif}(\mathcal{S}_{k})$. As noted in Appendix B, it suffices to prove the converse bound for the optimal decoder, that is, maximum likelihood estimation (MLE) given in (B.1). By symmetry with respect to re-ordering items, we can simply assume that the actual defective set is $s=[k]$, i.e., items $1,\dotsc,k$ are defective, and items $k+1,\dotsc,p$ are non-defective.
@@
For $j\in s$ and $j^{\prime}\in[p]\setminus s$ we will use the notation $\mathcal{M}_{j},\mathcal{M}_{j0},\mathcal{M}_{j1},\mathcal{N}_{0},\mathcal{N}_{00},\mathcal{N}_{01}$ and the corresponding cardinalities from Corollary B.1. Note that these sets and their cardinalities are deterministic given $\mathbf{X}_{s}$ and $\mathbf{Y}$. We consider $\ell=1$, and say that $(C,\zeta)\in(0,\infty)\times(0,1)$ is feasible if $\frac{Cn\nu e^{-\nu}}{k}$ and $\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}$ are integers. Given some feasible pair $(C,\zeta)$, we will consider $\mathcal{K}_{C,\zeta}=\{j\in s:M_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}\}$ with cardinality $k_{C,\zeta}=|\mathcal{K}_{C,\zeta}|$ as in (B.7). For a given pair $(j,j^{\prime})\in\mathcal{K}_{C,\zeta}\times([p]\setminus s)$, recall that $G_{j,j^{\prime},1}$ denotes the number of tests in $\mathcal{N}_{01}\cup\mathcal{M}_{j1}$ that contain the non-defective item $j^{\prime}$, and $G_{j,j^{\prime},2}$ denotes number of tests in $\mathcal{N}_{00}\cup\mathcal{M}_{j0}$ that contain the non-defective item $j^{\prime}$; see Corollary B.1.
We first invoke Corollary B.1 to identify sufficient conditions for the failure of MLE. In particular, MLE fails if for some feasible $(C,\zeta)\in(0,\infty)\times(0,1)$, the following two conditions (C1) and (C2) simultaneously hold:
@@
  • (C2) There exist $j\in\mathcal{K}_{C,\zeta}$ and $j^{\prime}\in\{k+1,\dotsc,p\}$ such that
    $G_{j,j^{\prime},1}-G_{j,j^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}.$ (C.1)
@@
We first derive a condition for ensuring $k_{C,\zeta}\geq 1$ with $1-o(1)$ probability. Fix a constant $\xi<1$ (to be chosen close to $1$ later) and suppose $k^{\xi}$ is an integer for notational convenience. For a given $j\in s$ and any test (say, with index $i$), we define the probability of test $i$ containing only one defective $j$ as
$P_{1}=\mathbb{P}(i\in\mathcal{M}_{j})=\frac{\nu}{k}\Big{(}1-\frac{\nu}{k}\Big{)}^{k-1}=\frac{\nu e^{-\nu}(1+o(1))}{k}.$ (C.2)
@@
$\|\bm{S}^{\prime}-\bm{T}\|_{\mathrm{TV}}=\left(\sum_{j=1}^{k}p_{j}\right)\left(\frac{1}{\sqrt{2\pi e}}+O\left(\frac{1}{\sqrt{N\sum_{j=1}^{k}p_{j}}}\right)\right).$ (C.3)
By Lemma C.1 and the fact that $n,k\to\infty$ as $p\to\infty$, we have
$d_{\mathrm{TV}}(\bm{M}_{\zeta},\hat{\bm{M}}_{\zeta})$ $=k^{\xi}P_{1}\Big{(}\frac{1}{\sqrt{2\pi e}}+O\Big{(}\frac{1}{\sqrt{nk^{\zeta}P_{1}}}\Big{)}\Big{)}$ (C.4)
@@
$k_{C,\zeta}$ $=|\mathcal{K}_{C,\zeta}|\geq|\mathcal{K}_{C,\zeta}\cap[k^{\xi}]|=\sum_{j=1}^{k^{\xi}}\mathbbm{1}\Big{(}M_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta Cn\nu e^{-\nu}}{k}\Big{)}.$ (C.5)
@@
$\mathbb{P}(k_{C,\zeta}=0)$ $\leq\mathbb{P}\left(\sum_{j=1}^{k^{\xi}}\mathbbm{1}\Big{(}M_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta Cn\nu e^{-\nu}}{k}\Big{)}=0\right)$ (C.6)
@@
$=\mathbb{P}\left(\sum_{j=1}^{k^{\xi}}\mathbbm{1}\Big{(}\hat{M}_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta Cn\nu e^{-\nu}}{k}\Big{)}=0\right)+o(1)$ (C.7)
$=\left[1-\mathbb{P}\left(\hat{M}_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta Cn\nu e^{-\nu}}{k}\right)\right]^{k^{\xi}}+o(1)$ (C.8)
$=\left[1-\mathbb{P}\left(\hat{M}_{j}=\frac{Cn\nu e^{-\nu}}{k}\right)\mathbb{P}\left(\text{Bin}\Big{(}\frac{Cn\nu e^{-\nu}}{k},\rho\Big{)}=\frac{\zeta Cn\nu e^{-\nu}}{k}\right)\right]^{k^{\xi}}+o(1),$ (C.9)
@@
where (C.8) holds because the $\hat{M}_{j}$ are independent, and the $\frac{Cn\nu e^{-\nu}}{k}$ tests in $\mathcal{M}_{j}$ are disjoint for different $j\in[k^{\xi}]$ (by the definition of $\mathcal{M}_{j}$), so the $k^{\xi}$ summands in (C.7) are independent; then, (C.9) holds because $\hat{M}_{j}$ is only related to the randomness of $\mathbf{X}_{s}$, while given $M_{j}=\frac{Cn\nu e^{-\nu}}{k}$, $M_{j0}\sim\text{Bin}(\frac{Cn\nu e^{-\nu}}{k},\rho)$ is only related to the randomness of $\mathbf{Z}$. Since $\hat{M}_{j}\sim\text{Poi}(nP_{1})$, we have
$\mathbb{P}\left(\hat{M}_{j}=\frac{Cn\nu e^{-\nu}}{k}\right)$ $=e^{-nP_{1}}\frac{(nP_{1})^{\frac{Cn\nu e^{-\nu}}{k}}}{\big{(}\frac{Cn\nu e^{-\nu}}{k}\big{)}!}$ (C.10)
$=\frac{\exp\big{(}-\frac{n\nu e^{-\nu}(1+o(1))}{k}\big{)}}{(C(1+o(1)))^{\frac{Cn\nu e^{-\nu}}{k}}}\cdot\frac{\big{(}\frac{Cn\nu e^{-\nu}}{k}\big{)}^{\frac{Cn\nu e^{-\nu}}{k}}}{\big{(}\frac{Cn\nu e^{-\nu}}{k}\big{)}!}$ (C.11)
$\geq\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{(}C\log C+1+o(1)\big{)}\right)\cdot\exp\left(\frac{Cn\nu e^{-\nu}}{k}(1+o(1))\right)$ (C.12)
$=\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{(}C\log C-C+1+o(1)\big{)}\right),$ (C.13)
@@
where (C.10) follows from (C.2), and (C.12) follows from Stirling’s approximation (specifically, $\frac{w^{w}}{w!}\geq\frac{e^{w-1}}{w}=e^{w(1+o(1))}$ for $w\to\infty$, with $w=\frac{Cn\nu e^{-\nu}}{k}$). Moreover, by anti-concentration ((A.4) in Lemma A.1), we have
$\mathbb{P}\left(\text{Bin}\Big{(}\frac{Cn\nu e^{-\nu}}{k},\rho\Big{)}=\frac{\zeta Cn\nu e^{-\nu}}{k}\right)$ $\geq\frac{1}{\sqrt{\frac{2Cn\nu e^{-\nu}}{k}}}\exp\left(-\frac{Cn\nu e^{-\nu}}{k}D(\zeta\|\rho)\right)$ (C.14)
$=\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{(}C\cdot D(\zeta\|\rho)+o(1)\big{)}\right),$ (C.15)
@@
where (C.15) holds because $\frac{n}{k}=\Theta(\log k)\to\infty$ allows us to incorporate the leading factor into the $o(1)$ term in the exponent; analogous simplifications will be used in the subsequent analysis. Combining the bounds in (C.13) and (C.15), we obtain
$\mathbb{P}\left(\hat{M}_{j}=\frac{Cn\nu e^{-\nu}}{k}\right)\mathbb{P}\left(\text{Bin}\Big{(}\frac{Cn\nu e^{-\nu}}{k},\rho\Big{)}=\frac{\zeta Cn\nu e^{-\nu}}{k}\right)$ (C.16)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \geq\exp\left(-\frac{n\nu e^{-\nu}}{k}\big{[}C\log C-C+C\cdot D(\zeta\|\rho)+1+o(1)\big{]}\right):=P_{2}.$ (C.17)
Substituting this into (C.9), we obtain
$\mathbb{P}(k_{C,\zeta}=0)$ $\leq(1-P_{2})^{k^{\xi}}+o(1)$ (C.18)
$\leq\exp\big{(}-k^{\xi}P_{2}\big{)}+o(1).$ (C.19)
@@
$f_{1}(C,\zeta,\rho):=C\log C-C+C\cdot D(\zeta\|\rho)+1,$ (C.20)
and noting that $\eta^{\prime}_{1}\log\frac{p}{k}\to\infty$ for any given $\eta^{\prime}_{1}>0$, the condition $k^{\xi}P_{2}\to\infty$ can be ensured by
$\xi\log k-\frac{n\nu e^{-\nu}}{k}\big{(}f_{1}(C,\zeta,\rho)+o(1)\big{)}\geq\eta^{\prime}_{1}\log\frac{p}{k}.$ (C.21)
Moreover, by $k=\Theta(p^{\theta})$ we have $\log k\sim\frac{\theta}{1-\theta}\log\frac{p}{k}$. Substituting this into (C.21), letting $\xi$ be arbitrarily close to $1$ and rearranging, we find that (C1) holds with $1-o(1)$ probability if the following holds for some $\eta_{1}>0$:
$n\leq\frac{(1-\eta_{1})\frac{\theta}{1-\theta}k\log\frac{p}{k}}{\nu e^{-\nu}\big{(}f_{1}(C,\zeta,\rho)+o(1)\big{)}}.$ (C.22)
@@
Given $k_{C,\zeta}\geq 1$ and fixing an arbitrary index $j\in\mathcal{K}_{C,\zeta}$, in this part we identify a sufficient condition for (C2). Most of our analysis will be conditioned on $(\mathbf{X}_{s},\mathbf{Y})$ and only based on the randomness of $\mathbf{X}_{[p]\setminus s}$, but we first deduce the high-probability behavior of the quantities $N_{01}=|\mathcal{N}_{01}|$ and $N_{00}=|\mathcal{N}_{00}|$ depending on randomness of $(\mathbf{X}_{s},\mathbf{Y})$. Specifically, $\mathcal{N}_{01}$ (resp., $\mathcal{N}_{00}$) consists of the positive (resp., negative) tests containing no defective, and since the probability of a specific test containing no defective is given by $(1-\frac{\nu}{k})^{k}=e^{-\nu}(1+o(1))$, we have $N_{01}\sim\text{Bin}(n,\rho e^{-\nu}(1+o(1)))$ and $N_{00}\sim\text{Bin}(n,(1-\rho)e^{-\nu}(1+o(1)))$. Then, by Hoeffding’s inequality, we have that
$\mathscr{A}_{1}:=\big{\{}N_{01}=\rho ne^{-\nu}(1+o(1)),\leavevmode\nobreak\ N_{00}=(1-\rho)ne^{-\nu}(1+o(1))\big{\}}$ (C.23)
@@
For the specific $j\in\mathcal{K}_{C,\zeta}$ we have $M_{j0}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}=o(n)$ and $M_{j1}=\frac{(1-\zeta)\cdot Cn\nu e^{-\nu}}{k}=o(n)$, and thus they are negligible compared to $N_{01},N_{00}$ specified in $\mathscr{A}_{1}$ in (C.23):
$|\mathcal{N}_{01}\cup\mathcal{M}_{j1}|=N_{01}+M_{j1}=\rho ne^{-\nu}(1+o(1)),$ (C.24)
$|\mathcal{N}_{00}\cup\mathcal{M}_{j0}|=N_{00}+M_{j0}=(1-\rho)ne^{-\nu}(1+o(1)).$ (C.25)
Note that these cardinalities correspond to counts of “relevant” tests with positive and negative outcomes respectively, and recall from Corollary B.1 that $G_{j,j^{\prime},1}$ and $G_{j,j^{\prime},2}$ count how many such tests include the non-defective $j^{\prime}$ (and thus impact the likelihood function of $(S\setminus\{j\})\cup\{j^{\prime}\}$ relative to $S$).